OpenStudy (thefluffmuncher):

Solve the system of equations. x + 3y = 1 –2x + y = –23 A. x = 22; y = 7 B. x = 16; y = –5 C. x = 10; y = –3 D. x = 14; y = 5

OpenStudy (anonymous):

It is C because it makes both true

OpenStudy (anonymous):

x+3y=1 10+3(-3)=1 1=1, so Thats correct Plug into the second equation: -2x+y=-23 -20+-3=-23 -23=-23

OpenStudy (thefluffmuncher):

thanks one more